3. (20 points) Algorithm Analysis and Recurrence There is a mystery function called Mystery(n) and the pseudocode of the algo pleas answer asap
Answer #1

From the current pseudo code ,we get the recurrence relation as -T(n) = 9T(n/3) + n

Masters Theorem for dividing function is as :

T(n) = aT(n/b) + f(n)  where f(n) is of form nklogp(n)

Solution for these equation by masters theorem is as

  1. if a > bk, then T(n) = θ(nlogba)
  2. if a = bk, then
    (a) if p > -1, then T(n) = θ(nlogba logp+1n)
    (b) if p = -1, then T(n) = θ(nlogba loglogn)
    (c) if p < -1, then T(n) = θ(nlogba)
  3. if a < bk, then
    (a) if p >= 0, then T(n) = θ(nk logpn)
    (b) if p < 0, then T(n) = θ(nk)

Hence for the current code ,we get the soln as :O(n2)